# Thermodynamic computing out of equilibrium

### v2

@@ I Introduction
While classical forms of computing view thermal fluctuations as an obstacle to computation Bennett (1982); Landauer (1991); Ceruzzi (2003), thermodynamic computing uses thermal fluctuations as a means of doing computation Fry (2005); Conte et al. (2019); Hylton (2020); Frank (2017); Wolpert (2019); Fry (2017); Conte et al. (2019); Wimsatt et al. (2021). Fluctuations can drive state changes in devices, and can be used to encode information. For instance, consider scalar degrees of freedom $x_{i}$ that interact via the bilinear couplings $J_{ij}x_{i}x_{j}$. If these degrees of freedom are placed in contact with a thermal bath at temperature $T$, then the equilibrium two-point correlations $\langle x_{i}x_{j}\rangle_{0}=k_{\rm B}T(J^{-1})_{ij}$ encode the elements of the matrix inverse of $J$. Thus measuring such correlations in equilibrium can be used to do matrix inversion Aifer et al. (2024); Melanson et al. (2023).
@@
A major focus of thermodynamic computing is to arrange for the equilibrium properties of a thermodynamic computer, described by the Boltzmann distribution, to correspond to the output of a specified computation. This approach is powerful, because knowing the interaction energy $U({\bm{x}})$ of the degrees of freedom ${\bm{x}}$ of a thermodynamic computer specifies the Boltzmann distribution $\rho_{0}({\bm{x}})=Z^{-1}{\rm e}^{-\beta U({\bm{x}})}$, where $\beta^{-1}\equiv k_{\rm B}T$ and $Z$ is the partition function. Thus by designing the interactions of the thermodynamic computer we can design its equilibrium properties Aifer et al. (2024). However, this approach comes with two challenges. One is that we need the computer to attain thermal equilibrium. In general, physical systems equilibrate on a broad range of timescales Arceri et al. (2022); Biroli and Garrahan (2013); Hagan et al. (2011); Whitelam and Jack (2015), and the equilibration times for even a simple thermodynamic computer can vary by orders of magnitude as its program is altered Whitelam (2024). A second challenge is that not every calculation can be represented by the Boltzmann distribution in an obvious way. For example, the matrix inversion described above can only be done if the matrix $J_{ij}$ is symmetric and positive definite. If it is not symmetric, we have $(J^{-1})_{ij}\neq(J^{-1})_{ji}$ in general, but we must have $\langle x_{i}x_{j}\rangle_{0}=\langle x_{j}x_{i}\rangle_{0}$, and so the relation described in the first paragraph cannot hold; if it is not positive definite, the system does not have a well-defined equilibrium distribution.
Figure 1: (a) A thermodynamic circuit whose interaction energy is given by Eq. (1) can function as a thermodynamic neuron. In panels (b–e) we show the equilibrium activation function $\langle x\rangle_{0}$ of the neuron, Eq. (2), as a function of the neuron input $I$, for the case $\beta=1$. The vector ${\bm{J}}=(J_{2},J_{3},J_{4})$ sets the values of the intrinsic couplings of the neuron. The top panels in (b–e) show the potential (1) at zero input, for the quadratic case (gray) and the case introduced in the lower panel (blue or green). (b) Purely quadratic (gray) and quartic (blue) cases. The purely quartic case (with $J_{4}>0$) is the simplest case that is thermodynamically stable and admits a nonlinear activation function. Both cases are shown for reference in the following panels. Inset: gradient of activation function on the same horizontal scale as the main panel. (c) Example in which a quadratic coupling is included with the quartic coupling. (d) Example in which a cubic coupling is included with the quartic coupling. (e) Example in which all three couplings are nonzero.
@@
We can sidestep these challenges by arranging for a thermodynamic computer to perform calculations out of equilibrium. Out of equilibrium we lose contact with the theoretical foundation provided by the Boltzmann distribution, and so we must find other ways of programming a thermodynamic computer in order to do specified calculations. Some exceptions to the equilibrium paradigm already exist. For instance, the matrix exponential $e^{-Jt}$ can be calculated at observation time $t$ Duffield et al. (2023), and nonequilibrium work measurements can be used to calculate the determinant of a matrix Aifer et al. (2024). However, no design exists for a general-purpose programmable thermodynamic computer that operates at specified observation times. Here we introduce such a design by introducing a thermodynamic computer that is analogous to a neural network. A thermodynamic computer of this nature is a nonlinear model that can serve as a universal function approximator, and can be programmed to perform arbitrary nonlinear computations at specified observation times. This is true whether or not the computer has attained thermodynamic equilibrium at those observation times.
@@
In more detail, we introduce a thermodynamic neuron, a fluctuating classical degree of freedom placed in contact with a heat bath and confined by a quartic potential  [^1] In general, the confining potential need not be exactly quartic, but it must be higher-order than quadratic and thermodynamically stable.. The equilibrium average of the neuron activation is a nonlinear function of the signal input to it, meaning that a network built from interacting thermodynamic neurons can function as a universal approximator Cybenko (1989); Hornik et al. (1989). Within a digital simulation of a thermodynamic computer we construct networks of such neurons using the bilinear interactions characteristic of existing thermodynamic computers Aifer et al. (2024); Melanson et al. (2023). We train these computers by genetic algorithm to perform nonlinear computations – expressing a nonlinear function and classifying MNIST – at specified times. Thermodynamic computers of this nature could be considered to be thermodynamic neural networks, or “thermoneural networks”.
@@
In Section II we introduce and analyze the properties of a simple thermodynamic circuit that can function as a thermodynamic neuron, a fluctuating degree of freedom whose output is a nonlinear function of its input. In Section III we simulate interacting networks of thermodynamic neurons, and train them by genetic algorithm to perform specified nonlinear computations. We conclude in Section IV.
II Thermodynamic neurons
II.1 Equilibrium behavior
To motivate the construction of a thermodynamic computer analogous to a neural network, consider the thermodynamic circuit shown in Fig. 1(a). This circuit represents a scalar degree of freedom $x$ that experiences the potential energy
$U_{\bm{J}}(x,I)=J_{2}x^{2}+J_{3}x^{3}+J_{4}x^{4}-Ix.$ (1)

### v3

# Nonlinear thermodynamic computing out of equilibrium

@@ Abstract
Introduction— In classical forms of computing, thermal fluctuations are an obstacle to computation Bennett (1982); Landauer (1991); Ceruzzi (2003); for thermodynamic computing, thermal fluctuations are the means of doing computation Fry (2005); Conte et al. (2019); Hylton (2020); Frank (2017); Wolpert (2019); Fry (2017); Conte et al. (2019); Wimsatt et al. (2021). Fluctuations can drive state changes in devices, and can be used to encode information. For instance, consider a thermodynamic computer comprising scalar degrees of freedom $x_{i}$ that interact via the bilinear couplings $J_{ij}x_{i}x_{j}$. If this computer is placed in contact with a thermal bath at temperature $T$, then its equilibrium two-point correlations $\langle x_{i}x_{j}\rangle_{0}=k_{\rm B}T(J^{-1})_{ij}$ encode the elements of the matrix inverse of $J$. Thus measuring such correlations in equilibrium can be used to do matrix inversion Aifer et al. (2024); Melanson et al. (2025).
@@
The current focus of thermodynamic computing is to arrange for the equilibrium properties of a thermodynamic computer, described by the Boltzmann distribution, to correspond to the output of a specified computation. This approach is powerful because the potential energy of a thermodynamic computer specifies the Boltzmann distribution, and so by designing the potential we can design the computer’s equilibrium properties Aifer et al. (2024). However, this approach comes with two challenges. One is that we need the computer to attain thermal equilibrium. In general, physical systems equilibrate on a broad range of timescales Arceri et al. (2022); Biroli and Garrahan (2013); Hagan et al. (2011); Whitelam and Jack (2015), and the equilibration times for even a simple thermodynamic computer can vary by orders of magnitude as its program is altered Whitelam (2025a). A second challenge is that not every calculation can be represented by the Boltzmann distribution in an obvious way. For example, the matrix inversion described above can only be done if the matrix $J_{ij}$ is positive definite; if not, the system does not possess a well-defined equilibrium distribution.
Figure 1: (a) A thermodynamic circuit whose interaction energy is given by Eq. (1) can function as a thermodynamic neuron. The circle represents a scalar degree of freedom $x$. The curved line represents its intrinsic energy, the terms in ${\bm{J}}=(J_{2},J_{3},J_{4})$ in Eq. (1). The straight line represent an input signal or bias, the term in II in Eq. (1). (b) Equilibrium activation function $\langle x\rangle_{0}$ of the neuron, Eq. (2), as a function of the neuron input $I$, for the case $\beta=1$. The vector ${\bm{J}}=(J_{2},J_{3},J_{4})$ sets the values of the intrinsic couplings of the neuron. The quadratic-potential activation function is linear, while the quartic-potential activation function is nonlinear. (c) Dynamical evolution (3) of the quadratic-potential neuron, for $\beta=100$, for 11 evenly-spaced values of $I$. (d) The same for the quartic-potential neuron. For times longer than some short threshold, the finite-time response is a nonlinear function of $I$.
@@
We can sidestep these challenges by arranging for a thermodynamic computer to perform calculations out of equilibrium. Out of equilibrium we lose contact with the theoretical foundation provided by the Boltzmann distribution, and so we must find other ways of programming a thermodynamic computer in order to do specified calculations. Some exceptions to the equilibrium paradigm already exist. For instance, the matrix exponential $e^{-Jt}$ can be calculated at observation time $t$ Duffield et al. (2023), and nonequilibrium work measurements can be used to calculate the determinant of a matrix Aifer et al. (2024). However, no design exists for a thermodynamic computer that operates at specified observation times and is programmable in a general sense, meaning that it is capable of approximating arbitrary continuous functions. Here we provide such a design by introducing a thermodynamic computer that is analogous to a neural network. A thermodynamic computer of this nature is a nonlinear model that can serve as a universal function approximator, and can be programmed to perform arbitrary nonlinear computations at specified observation times. This is true whether or not the computer has attained thermodynamic equilibrium at those observation times.
The equilibrium thermodynamic computers of  Refs. Aifer et al. (2024); Melanson et al. (2025), which motivate this work, can be viewed as continuous-spin analogs of Hopfield networks Hopfield (1982) or Boltzmann machines Hinton (2017); Salakhutdinov and Hinton (2009), statistical mechanical models that represent probability distributions over binary variables. Since a Boltzmann machine encodes information in its equilibrium or Boltzmann distribution, one could refer to the models of  Refs. Aifer et al. (2024); Melanson et al. (2025) as Boltzmann computers: they are designed to sample from the Boltzmann equilibrium corresponding to the system’s potential energy function. In such models, the specific form of the dynamics, whether overdamped or underdamped, is less important than the fact that it is microscopically reversible and converges to the desired equilibrium.
Our contribution is to explore the computational capabilities of such systems when extended to include nonlinear neuron potentials and operated out of equilibrium. Specifically, we encode the outcome of a computation in the dynamical trajectories of the system, without appealing to the Boltzmann distribution. A computer operating in this mode, using Langevin trajectories to perform a computation, could be termed a Langevin computer.
The theoretical framework describing thermodynamic computing includes nonequilibrium statistical mechanics, stochastic thermodynamics, and information theory. These fields offer a set of tools for analyzing small, fluctuating systems, such as fluctuation theorems, thermodynamic speed limits, and uncertainty relations Seifert (2012); Wolpert et al. (2024); Sagawa (2014). For example, the Jarzynski equality, combined with nonequilibrium work measurements, can be used to compute the determinant of a matrix Aifer et al. (2024). Thermodynamic computers often use added noise to accelerate computational, and the energetic cost of this addition can be quantified within stochastic thermodynamics Whitelam (2025a). Learning rules can be linked to thermodynamic quantities: for instance, a generative thermodynamic computer trained by gradient descent minimizes heat dissipation during training Whitelam (2025b).
Thermodynamic neurons — In more detail, we introduce the thermodynamic circuit shown in Fig. 1(a), a fluctuating classical degree of freedom placed in contact with a heat bath and confined by a quartic potential  [^1] In general, the confining potential need not be exactly quartic, but it must be higher-order than quadratic and thermodynamically stable (i.e. bounded from below). We can consider Eq. (1) to represent a Maclaurin expansion, in powers of $x$, of an arbitrary nonlinear thermodynamic circuit.. This circuit represents a scalar degree of freedom $x$ that experiences the potential energy
$U_{\bm{J}}(x,I)=J_{2}x^{2}+J_{3}x^{3}+J_{4}x^{4}-Ix.$ (1)
Let the neuron be put in contact with a thermal bath at temperature $T$. In thermal equilibrium, the output of the neuron has the mean value
$\langle x\rangle_{0}\equiv\frac{\int{\rm d}x\,x\,{\rm e}^{-\beta U_{\bm{J}}(x,I)}}{\int{\rm d}x\,{\rm e}^{-\beta U_{\bm{J}}(x,I)}}.$ (2)
When the neuron potential is purely quadratic, i.e. ${\bm{J}}=(J_{2},0,0)$, the integrals in (2) can be solved analytically, giving the linear form $m=I/(2J_{2})$. This form is plotted, for $J_{2}=1$, as a gray line in Fig. 1(b); the horizontal dotted black line denotes the value zero. In this case the equilibrium activation function of the neuron is linear, meaning that networks of such neurons in equilibrium cannot serve as universal function approximators Cybenko (1989); Hornik et al. (1989). The simplest case that is thermodynamically stable (i.e. the potential is bounded from below) and admits a nonlinear equilibrium activation function is the purely quartic case, ${\bm{J}}=(0,0,J_{4})$, with $J_{4}>0$. This case is shown as a blue line in Fig. 1(b), with $J_{4}=1$  [^2] In the purely quartic case the equilibrium activation function can be expressed analytically, in terms of the hypergeometric function.. The equilibrium activation function is nonlinear: its gradient (shown inset) is largest near the origin, and decreases as $|I|$ becomes large.
An additional design consideration is the variance $\sigma^{2}=\langle x^{2}\rangle_{0}-\langle x\rangle_{0}^{2}$ of the neuron’s equilibrium fluctuations. We show in the Supplementary Material that adding a quadratic term to the quartic term suppresses the neuron’s fluctuations in equilibrium. The larger the fluctuations of a thermodynamic neuron’s output, the more samples will be required to compute a meaningful signal when observing a computer built from such neurons. For these reasons we choose our default neuron parameters to be ${\bm{J}}=(1,0,1)$: the quartic coupling induces nonlinearity, while the quadratic coupling serves to suppress fluctuations near $I=0$.
The nonequilibrium response of the quartic-potential thermodynamic neuron is also, in general, nonlinear with input. Thermodynamic computers operate under Langevin dynamics Aifer et al. (2024); Melanson et al. (2025), and here we consider the overdamped dynamics
$\dot{x}_{i}=-\mu\frac{\partial V({\bm{x}},{\bm{I}})}{\partial x_{i}}+\sqrt{2\mu k_{\rm B}T}\,\eta_{i}(t),$ (3)
where $i$ labels the neuron; $V({\bm{x}},{\bm{I}})$ is the computer’s potential given input $\bm{I}$}; and the Gaussian white noise terms satisfy $\langle\eta_{i}(t)\rangle=0$ and $\langle\eta_{i}(t)\eta_{j}(t^{\prime})\rangle=\delta_{ij}\delta(t-t^{\prime})$. The mobility parameter $\mu$ sets the basic time constant of the computer. For the thermodynamic computers of Refs. Aifer et al. (2024); Melanson et al. (2025), $\mu^{-1}\sim 1$ microsecond. For damped oscillators made from mechanical elements Dago et al. (2021) or Josephson junctions Ray and Crutchfield (2023), $\mu^{-1}$ is of order a millisecond or a nanosecond, respectively.
In Fig. 1(c,d) we show the dynamical evolution of a single neuron $x$ in the potential $V({\bm{x}},{\bm{I}})=U_{\bm{J}}(x,I)$, under the dynamics (3), starting from $x=0$. Panel (c) shows the case of the quadratic potential ${\bm{J}}=(1,0,0)$, for various fixed values of the input $I$. The neuron is initially out of equilibrium, converging to equilibrium in about 2 time units (time is expressed in units of $\mu^{-1}$). As described by Fig. 1(b), the mean value of $x$ in equilibrium is a linear function of $I$.
Panel (d) shows the case of a quartic potential ${\bm{J}}=(0,0,1)$. For observation times larger than some threshold, the neuron’s output is a nonlinear function $x(I)$, and eventually converges to the equilibrium activation function $\langle x(I)\rangle_{0}$, which we have designed to be nonlinear. A network of such neurons, observed on similar timescales, can therefore function as a universal approximator. (In a related vein, nonlinear physical neural networks that operate on the energy scales of classical computing have recently been trained to do nonlinear computations Dillavou et al. (2024).)
Figure 2: (a,b) Elements of a thermodynamic computer analogous to a neural network. (a) The thermodynamic neurons described in Fig. 1 are connected by bilinear couplings. Top: a single circle implies a neuron $x_{i}$ of the type described in Fig. 1, with an input (bias) $I=b_{i}$ and a nonlinear potential parameterized by ${\bm{J}}=(J_{1},J_{2},J_{3})$. Bottom: lines between circles imply a bilinear coupling $J_{ij}x_{i}x_{j}$. (b) With the visual shorthand described in panel (a), we consider layered networks of such neurons, with adjacent layers coupled all-to-all, having total potential energy (4). (c) Training a simulation model of a thermodynamic computer to express a nonlinear function at a specified observation time. We show loss (8) as a function of evolutionary time $n$ for a layered thermodynamic computer built from quadratic neurons (gray) or quadratic-quartic neurons (green). (d) Output (9) at observation time $t_{\rm f}=1$ of the linear computer (gray) and the nonlinear computer (green), as a function of the input $z$, averaged over $M=10^{3}$ samples. The target function is shown as a black line. (e) Mean neuron activations measured at observation time $t_{\rm f}$ as a function of the neuron inputs at the same time, for the nonlinear model. The color band denotes $\pm$ one standard deviation. (f) Output (9) at time $t_{\rm f}=1$ of the trained nonlinear thermodynamic computer as a function of input $z$, computed using $M$ samples. The target function is shown as a black line (training was done using $M=10^{3}$ samples). (g) Output (9) at various observation times $t_{\rm f}$ of the trained nonlinear thermodynamic computer, as a function of input $z$, computed using $M=10^{3}$ samples. The target function is shown as a black line. The computer is trained so that it reproduces the target function when observed at time $t_{\rm f}=1$.
Thermodynamic neural network — To illustrate this statement, consider a thermodynamic computer consisting of a network of $N$ thermodynamic neurons $x_{i}$, with potential energy function
$V({\bm{x}},{\bm{I}})=V_{\rm int}({\bm{x}})+V_{\rm ext}({\bm{x}},{\bm{I}})$ (4)
Here
$V_{\rm int}({\bm{x}})=\sum_{i=1}^{N}U_{\bm{J}}(x_{i},b_{i})+\sum_{(ij)}J_{ij}x_{i}x_{j}.$ (5)
is the internal potential of the computer, accounting for neuron self-interactions and connections between neurons. The first sum in (5) runs over $N$ single-neuron energy terms (1), while the second sum runs over all distinct pairs of connected neurons. We use the bilinear interaction of Refs. Aifer et al. (2024); Melanson et al. (2025). The computers described in those papers use an all-to-all coupling; here, to make contact with existing neural-network designs, we consider the layered structure shown in Fig. 2, with all-to-all connections between layers. This design mimics that of a conventional deep fully-connected neural network. However, unlike in a conventional deep neural network, in which information flows from the input layer to the output layer, the bilinear interaction $J_{ij}x_{i}x_{j}$ ensures that neuron $i$ communicates with neuron $j$, and vice versa, and so information flows forward and backward between the layers of the thermodynamic computer.
$V_{\rm ext}({\bm{x}},{\bm{I}})=\sum_{{\rm inputs}(ij)}W_{ij}I_{i}x_{j},$ (6)
where the sum runs over all connections between the external inputs $I_{i}$ and the input neurons $x_{j}$ (here the top-layer neurons), mediated by the parameters $W_{ij}$.
The computer evolves according to the Langevin dynamics (3). Because the computer is noisy, we wish to take $M$ samples of its output and average over these samples. In the main text we do this by reset-sampling: starting from zero neuron activation ${\bm{x}}={\bm{0}}$, we run the computer for time $t_{\rm f}$, observe the outcome, reset the neuron activations to zero, and repeat the procedure, gathering $M$ samples in total (in the Supplementary Material we also consider the case of serial sampling). Reset sampling naturally lends itself to parallelization: the $M$ samples can be computed independently, on distinct copies of the thermodynamic computer if such copies are available. Reset sampling can also be done using a single computer whose neurons are reset periodically.
We designate the final-layer neurons of the computer as its outputs, and calculate the reset-sampling averages
where the subscript ‘r’ indicates ‘reset-sampling’, and the sum runs over $M$ independent trajectories $\alpha$ of the dynamics (3). The only requirement on $t_{\rm f}$ is that it is long enough that the effective activation function of the neuron is nonlinear. We set $t_{\rm f}=1$, in units of $\mu^{-1}$.
Programming a thermodynamic computer— A nonlinear thermodynamic computer can be programmed to perform arbitrary nonlinear computations at specified times. We first consider the task of expressing a nonlinear function of a single variable. In this case the computer has one input, $I=z$. We define the target function $y_{0}(z)\equiv\cos(2\pi z)$ and the loss function
$\phi\equiv K^{-1}\sum_{j=1}^{K}\left(y_{0}(z_{j})-y(z_{j})\right)^{2},$ (8)
where the sum runs over $K=250$ evenly-spaced points $z_{j}=j/(K-1)$ on the interval $z\in[0,1]$. The quantity
is the output of the thermodynamic computer, given the input $z$, averaged over $M=10^{3}$ samples. The subscript ‘r,s’ indicates that averages are taken either in reset-sampling or in serial-sampling mode. If the former, each of the $M$ samples is obtained from an independent trajectory. If the latter, all $M$ samples are obtained from a single trajectory.
The adjustable parameters of the computer are ${\bm{\theta}}=\{W_{ij}\}\cup\{b_{i}\}\cup\{J_{ij}\}\cup\{f_{i}\}$. Here $\{W_{ij}\}$ is the set of input weights specified by Eq. (6); $\{b_{i}\}$ is the set of biases specified by Eq. (5); $\{J_{ij}\}$ is the set of connections specified by the same equation; and $\{f_{i}\}$ is a set of weights that couple to the output neurons. To program the computer we adjust the parameters $\bm{\theta}$} using a genetic algorithm instructed to minimize a loss function $\phi$, using an efficient GPU implementation (see Supplementary Material).
Following training, the identity of the computer’s parameters are fixed, and the computer can be run for any chosen input. The parameters of the digital model of the thermodynamic computer could in principle be implemented in hardware, resulting in a device designed to output a specified computation at a specified time, powered by thermal fluctuations. We note that if the hardware implementation is not an exact copy of the digital model, the genetic-algorithm training could be continued directly in hardware: the procedure can be applied to an experimental system exactly as it is applied to a simulation model Sabattini et al. (2024).
We choose a layered computer design of width 8 and depth 4. We consider two types of thermodynamic neuron: a quadratic-quartic thermodynamic neuron, ${\bm{J}}=(1,0,1)$, which gives rise to a nonlinear computer, and a quadratic thermodynamic neuron, ${\bm{J}}=(1,0,0)$, which gives rise to a linear computer. We take $\beta=10$, so that the neuron energy scale is 10 times that of the thermal energy.
In Fig. 2(c) we show the loss as a function of evolutionary time for the two models. The linear model fails to train – it cannot express a nonlinear function of the input variable – while the nonlinear model learns steadily, reaching a small value of the loss. Panel (d) shows the output functions learned by the two models: the nonlinear model has learned a good approximation of the target cosine function. The intrinsic noise of the computer is visible in the output, but for $M=10^{3}$ samples, for each value of $z$, the mean output signal of the computer exceeds the scale of the noise by a considerable margin.
Panel (e) of Fig. 2 shows the sampled neuron outputs as a function of the neuron inputs (the inputs being all signals in to the neuron, excepting the thermal noise) at the designated observation time. The nonlinear model possesses a nonlinear finite-time activation function, explaining the computer’s ability to learn an arbitrary nonlinear function. By contrast, the quadratic-neuron computer is at all times a linear model, and is unable to express a nonlinear function.
Training is done using $M=10^{3}$ samples for each value of the input $z$, but the trained computer can be used with fewer samples if desired. In Fig. 2(f) we show the output of the trained thermodynamic computer, as a function of the input $z$, for a range of values of $M$.
Training in reset-sampling mode results in a thermodynamic computer programmed to express the target function at a prescribed observation time $t_{\rm f}=1$. In Fig. 2(g) we show the output of the computer at a range of observation times. The output of the computer varies as a function of time, and is equal to the target function only at the prescribed observation time. The output of the computer in equilibrium (corresponding to the long-time limit) is considerably different to the target function. In this example, therefore, the programmed thermodynamic computer operates far from equilibrium.
Machine learning with a thermodynamic computer — Having confirmed the ability of a network of nonlinear thermodynamic neurons to express an arbitrary nonlinear function, we now consider a standard benchmark in machine learning, classifying the MNIST data set LeCun et al. (1998a). MNIST consists of greyscale images of $70,000$ handwritten digits on a grid of $28\times 28$ pixels, each digit belonging to one of ten classes $C\in[0,9]$.
Figure 3: Training a simulation model of a thermodynamic computer to classify MNIST. The computer, which consists of a 3-layer network of quadratic-quartic $(1,0,1)$ neurons, is trained in reset-sampling mode, using $M=10^{3}$ samples taken at observation time $t_{\rm f}=1$. (a) Loss (cross-entropy) and training-set classification accuracy as a function of evolutionary time $n$. (b) For a single digit, an 8, we show the probability distribution, taken over $10^{5}$ samples, of the computer’s per-sample class score. The mean value of each distribution, which is the value used for classification, is indicated at the top of the panel. The correct distribution is shown in green, the others in shades of blue. (c) The class probabilities of the computer, upon being shown the indicated digit, for various observation time $t_{\rm f}$. The computer is trained to classify the digit at an observation time $t_{\rm f}=1$ (vertical dotted line). (d) Test-set classification accuracy of the trained computer, as a function of the number of samples $M$ generated by the computer (each taken at observation time $t_{\rm f}=1$). (e) Test-set accuracy of the trained computer at a range of energy scales. The computer is trained at the energy scale $J=10k_{\rm B}T$.
We simulate a 3-layer thermodynamic computer with quadratic-quartic $(1,0,1)$ neurons. Each layer has 32 neurons. Each neuron in the input layer couples to all the pixels $I_{i}$ of an MNIST digit via Eq. (6). The output layer of 32 neurons is used to construct the computer’s prediction for the class of MNIST digit ${\bm{I}}_{j}$, and we train the computer by genetic algorithm, instructing it to minimize the cross-entropy between the class probabilities predicted by the thermodynamic computer and the ground-truth labels (see Supplementary Material).
In Fig. 3(a) we show the loss as a function of evolutionary time $n$ as the computer is trained to classify MNIST. The computer learns steadily under the action of the genetic algorithm. Panel (b) shows the corresponding training-set classification accuracy (which can be observed but is not used during training). The corresponding test-set accuracy after training is about 93%, which is not state-of-the-art – many other methods classify MNIST with greater accuracy LeCun et al. (1998b) – but it is more accurate than a linear classifier, and confirms the ability of a thermodynamic computer to address standard machine-learning problems. As with conventional neural networks, better accuracy will be achieved using different computer designs and methods of training. Here our aim is to show proof of principle: if implemented in hardware, this thermodynamic computer would be able, powered only by thermal fluctuations, to classify MNIST digits.
In Fig. 3(b) we show the output of the trained computer when presented with a single digit, an 8, which it correctly classifies. We distinguish between the class score $Y^{(C)}$, the raw output of the thermodynamic computer prior to normalization, and the class probability $p^{(C)}$, which results from applying a softmax transformation to averaged score. We use the term class prediction to refer to the class with the highest score, i.e., the computer’s predicted label for the digit. We plot the probability distribution, taken over $10^{5}$ samples, of the computer’s class score $Y^{(C)}$ for the digit. The mean value of each distribution (the value used for classification) is indicated at the top of the panel. The correct distribution is shown in green, with the others in shades of blue. In panel (c) we plot the value of the trained computer’s 10 class probabilities upon being shown the indicated digit. The computer is trained to classify the digit at an observation time $t_{\rm f}=1$. In this case the computer has attained a steady-state dynamics at the specified observation time, but this is not a general phenomenon: when presented with other digits, the computer’s neurons are still evolving at $t_{\rm f}=1$, and so in general the computer operates out of equilibrium.
Fig. 3(d) shows the test-set accuracy of the computer upon collecting $M$ samples. Training was done with $M=10^{3}$ samples, but similar accuracy can be obtained using only about 20 samples. Thus if implemented in hardware, the thermodynamic computer could perform classification relatively efficiently.
The ability of a trained computer to operate at different noise scales depends on the nature of the problem it is trained for and its architecture. In Fig. 3(e) we show the test-set accuracy of the trained thermodynamic computer at a range of energy scales, calculated by holding fixed the parameters of the thermodynamic computer and varying the noise strength (see Supplementary Material). The computer, trained at the energy scale $J=10k_{\rm B}T$, performs essentially as well when subjected to noise comparable to its own energy scale ($J=k_{\rm B}T$). This result shows the ability of the computer to operate reliably throughout the regime characteristic of thermodynamic computing $(k_{\rm B}T\lesssim J\ll 10^{3}k_{\rm B}T)$, and shows its output to be robust to small changes in noise level (which might occur if a device becomes hot during computation). For sufficiently large noise levels the computer’s performance begins to decline.
Conclusions— Classical computing aims to suppress thermal fluctuations, while thermodynamic computing uses them. Most existing thermodynamic computing focus on doing linear algebra in equilibrium, using the Boltzmann distribution to encode computations Aifer et al. (2024); Melanson et al. (2025). A thermodynamic computer used in this way could be described as a Boltzmann computer. However, equilibration can be slow, and not all problems can be encoded as a Boltzmann distribution. This paper presents an alternative approach, in which a nonlinear thermodynamic computer is trained to perform nonlinear calculations at specified times. A thermodynamic computer trained to perform computations using Langevin trajectories could be called a Langevin computer.
The core component of the design, a thermodynamic neuron, is a fluctuating degree of freedom confined by a non-quadratic potential, allowing networks of such neurons to act as universal approximators. A recent paper Lipka-Bartosik et al. (2024) presented the design for a thermodynamic neuron realized by qubits coupled to multiple thermal baths. The design presented here is based on a fully classical model of a nonlinear thermodynamic neuron, and requires only a single thermal bath. Our proposed design, sketched in Fig. 2, could be implemented using existing hardware elements: the neurons could be realized by RLC Melanson et al. (2025) circuits with nonlinear components, or by superconducting circuits with Josephson junctions Ray and Crutchfield (2023). A thermodynamic computer of this nature could be programmed to perform arbitrary nonlinear computations at specified times, powered by thermal fluctuations.
Acknowledgments— This work was done at the Molecular Foundry, supported by the Office of Science, Office of Basic Energy Sciences, of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231. C.C. was supported by a Francqui Fellowship of the Belgian American Educational Foundation, and by the US DOE Office of Science Scientific User Facilities AI/ML project “A digital twin for spatiotemporally resolved experiments”.
Supplementary Material
S1 Overview
@@
In this supplement we describe the design and training of the nonlinear thermodynamic computer presented in the main text. The core of the design is a thermodynamic neuron, a fluctuating classical degree of freedom placed in contact with a heat bath and confined by a quartic potential  [^3] In general, the confining potential need not be exactly quartic, but it must be higher-order than quadratic and thermodynamically stable.. The equilibrium average of the neuron activation is a nonlinear function of the signal input to it, meaning that a network built from interacting thermodynamic neurons can function as a universal approximator Cybenko (1989); Hornik et al. (1989). Within a digital simulation of a thermodynamic computer we construct networks of such neurons using the bilinear interactions characteristic of existing thermodynamic computers Aifer et al. (2024); Melanson et al. (2025). We train these computers by genetic algorithm to perform nonlinear computations – expressing a nonlinear function and classifying MNIST – at specified times. Thermodynamic computers of this nature could be considered to be thermodynamic neural networks, or “thermoneural networks”.
@@
In Section S2 we introduce and analyze the properties of a simple thermodynamic circuit that can function as a thermodynamic neuron, a fluctuating degree of freedom whose output is a nonlinear function of its input. In Section S3 we simulate interacting networks of thermodynamic neurons, and train them by genetic algorithm to perform specified nonlinear computations.
S2 Thermodynamic neurons
Figure S1: (a) A thermodynamic circuit whose interaction energy is given by Eq. (S1) can function as a thermodynamic neuron. In panels (b–e) we show the equilibrium activation function $\langle x\rangle_{0}$ of the neuron, Eq. (S2), as a function of the neuron input $I$, for the case $\beta=1$. The vector ${\bm{J}}=(J_{2},J_{3},J_{4})$ sets the values of the intrinsic couplings of the neuron. The top panels in (b–e) show the potential (S1) at zero input, for the quadratic case (gray) and the case introduced in the lower panel (blue or green). (b) Purely quadratic (gray) and quartic (blue) cases. The purely quartic case (with $J_{4}>0$) is the simplest case that is thermodynamically stable and admits a nonlinear activation function. Both cases are shown for reference in the following panels. Inset: gradient of activation function on the same horizontal scale as the main panel. (c) Example in which a quadratic coupling is included with the quartic coupling. (d) Example in which a cubic coupling is included with the quartic coupling. (e) Example in which all three couplings are nonzero.
S2.1 Equilibrium behavior
To motivate the construction of a thermodynamic computer analogous to a neural network, consider the thermodynamic circuit shown in Fig. S1(a). This circuit represents a scalar degree of freedom $x$ that experiences the potential energy
$U_{\bm{J}}(x,I)=J_{2}x^{2}+J_{3}x^{3}+J_{4}x^{4}-Ix.$ (S1)